Operations Report – January 24th

SOL: 12

Name of person filing report: Freddy Castaneda

Non-nominal systems: Nothing new to report.

Notes on non-nominal systems: None

Generator (hours run): 7hr 0min; Turned on last night (23Jan2019) at 16:30; Stopped running during the night, about 23:30. It was turned off all night. This morning the contractor checked the generator, it will be turned off for three days.

Solar SOC – Turned on (23Jan2019) 84%; Turned off (24Jan2019) 95%; Turned on (24Jan2019) 86%

Diesel Reading – Under 1/2 tank.

Propane Reading – Still 80%

Ethanol Free Gasoline – Nothing to report.

Water (auxiliary tank) – Not in use.

Water (static tank) – Just over 20%; 115 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – About 37%; 110 gallons

Water (loft) – At level marker 12

Static to Loft Pump used – Yes; At 18:05 to refill tank

Water Meter: 01404464

Toilet tank emptied: No.

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Not used.

Hours: 70.3

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Opportunity rover used: Not used.

Hours: 46.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Not used.

Hours: 72.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Brakes on Opportunity still not working properly.

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: None

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? Not used.

CrewCar used and why, where? Off campus

General notes and comments: Nothing to report.

Summary of internet: Nothing to report.

Summary of suits and radios: Suit #3 without fuse, Suit #10 Pending to check battery functionality on field, Suit #11 Pending to check battery functionality on field.

Summary of Hab operations: Vacuum filter and tissue paper received.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: Nothing to report.